OpenStudy (anonymous):

Neveah has created the function f(x) = the quantity of 2x plus 4, divided by 3 to represent the growth of her hair, where x represents the number of months since her last hair cut. Neveah discovers that using the inverse function to solve for x = 6, she can predict when her hair will have grown 6 inches. Explain to Neveah how to accomplish this, using complete sentences.

OpenStudy (anonymous):

@marigirl

OpenStudy (anonymous):

Not looking for sentences, just the idea of it

OpenStudy (anonymous):

@heyyeahyou

OpenStudy (marigirl):

do u know how to work out an inverse function for this equation y= (2x+4)/3

OpenStudy (marigirl):

basically what you do is make x the subject y= (2x+4)/3 , then switch the x and y around.. (see here for more info: http://www.purplemath.com/modules/invrsfcn3.htm)

OpenStudy (marigirl):

so if we apply it to here y= (2x+4)/3 the inverse will be y=(3x-4)/2 (PLEASE let me know if u want me to go through step by step)

OpenStudy (anonymous):

I would like you to go step by step only if you have time so that I can understand it to write the sentences.

OpenStudy (marigirl):

yes, we need to work out the inverse function for y= (2x+4)/3 now solve for 'x=..' y= (2x+4)/3 3y=2x+4 3y-4=2x (3y-4)/2=x now that you have "x =", I'll switch x and y; the "y =" is the inverse. y=(3x-4)/2

OpenStudy (marigirl):

now lets solve for x=6 y=(3x-4)/2 y=(3(6)-4/2 y=7

OpenStudy (marigirl):

i hope that i make sense...

OpenStudy (anonymous):

That all made perfect sense! Is that the end of the problem? That was a great explanation

OpenStudy (marigirl):

Neveah discovers that using the inverse function to solve for x = 6, she can predict when her hair will have grown 6 inches..... I think I am interpreting this correctly inverse funtion y=(3x-4)/(2) then when we solve for x=6 we got y=7 so this means at 7 months her hair would have grown 6 inches.... i hope im right

OpenStudy (anonymous):

Thank you so much again! I believe that you are right. :)